OpenStudy (anonymous):

find the general solution of the equation y'=(x^2+3y^2)/(2xy)

OpenStudy (anonymous):

i get y = e^(3x^2/8)

OpenStudy (anonymous):

can anyone verify this?

OpenStudy (anonymous):

wolfram disagrees.

OpenStudy (anonymous):

It would take me too long to look up how to do it though I'm afraid. I need to review diff-eqs.

myininaya (myininaya):

\[y'=\frac{x^2+3y^2}{2xy}\] \[2xy y'=x^2+3y^2\] \[2xyy'-3y^2=x^2\] \[x \cdot 2yy'-3y^2=x^2\] ------------------------ Recall (y^2)'=2yy' \[2yy'-3\frac{1}{x} y^2=x\] assuming \[x \neq 0\] now we need to multiply v on both sides such that we can write \[(vy^2)'\] on the right hand side so how do we choose this v well let's think how can we get v 2yy'+v' y^2 on right hand side we need v'=-3/x so v is a fuction of x so we can write \[\frac{dv}{dx}=\frac{-3}{x}\] we can use separation of variables to find what v needs to be \[dv= \frac{-3}{x} dx\] integrate both sides \[v=-3 \ln|x|+C_1\] let's just use 0 for C_1 so we have \[v=-3 \ln|x|\] so we have \[-3 \ln|x| 2y y'+ \frac{-3}{x} y^2 =-3x \ln|x|\] so cleaning this up a bit we have \[-3 \ln|x| (y^2)'+(-3 \ln|x|)'y^2=-3x \ln|x|\] so now we can write \[(-3 \ln|x| y^2)'=-3 x \ln|x|\] now integrate both sides \[-3 \ln|x| y^2+K_1=-3 (\frac{x^2}{2} \ln|x|-\int\limits_{}^{} \frac{x^2}{2} \frac{1}{x} dx +K_2)\] \[-3 \ln|x| y^2+K_1=-3(\frac{x^2}{2} \ln|x|-\frac{1}{2} \cdot \frac{x^2}{2}+K_2)\] Note: \[-3K_2-K_1=K\] \[-3\ln|x|y^2=\frac{-3x^2}{x}\ln|x|+\frac{3x^2}{4}+K\]

myininaya (myininaya):

does wolfram agree with that?

OpenStudy (anonymous):

wolfram just said it was x sqrt (x-1) + c

OpenStudy (anonymous):

but i have no clue how to get anywhere near that

myininaya (myininaya):

we can see what happens if we tried so solve for y

myininaya (myininaya):

oops that was a mistake on the last line that x is suppose to be a 2 on the bottom of that one fraction

myininaya (myininaya):

\[-3\ln|x|y^2=\frac{-3x^2}{2} \ln|x|+\frac{3x^2}{4}+K\]

myininaya (myininaya):

\[y^2=\frac{x^2}{2}+\frac{-x^2}{4 \ln|x|}+\frac{K}{-3 \ln|x|}\] hmmm.... i don't this will end up looking like what they got

myininaya (myininaya):

:(

myininaya (myininaya):

i did make a mistake i will redo ok?

OpenStudy (anonymous):

i owe you a huge favor for this one myinin

myininaya (myininaya):

:)

myininaya (myininaya):

no favors needed